[Jamie, Koji, Suresh]
We replaced splitters in several demod boards as per the table given below:
While doing a rough check of the boards I noticed that the REFL11 demod board had no signal on the Q output.
Rana also advised that we must use the boards which have the piggy-back amplifiers on those signals which are most useful. We referred to Alberto's thesis and chose POY55 (MICH and SRCL), REFL11(PRCL) and AS55 (DARM) as the most useful signals. We currently have these amps on AS11, REFL11 and AS55. We need to convert either AS11 or REFL11 into a POY55. Since we need to troubleshoot REFL11, I thought we might as well modify that and in the process also fix its Q output. So I renamed AS11 as REFL11 and will convert the old REFL11 into POY55 tomorrow.
Power splitter of different types have different pin-outs. The way we mount a splitter depends on which type we are using. I will detail the mounting scheme in a separate elog tomorrow.
The PSL temperature box has returned to service, with some circuit modifications. The 1k resistors on all the temp. sensor inputs (R3, R4, R7, R8, R12, R12) were changed to 0 Ohm. Also, the 10k resistors R26, R28, R29, and R30 were changed to 10.2k metal film. The DCC document will be updated shortly. There is now an offset in the MINCOMEAS channel compared to the others, which will be corrected in the morning after looking at the overnight trend.
This is what was done in past two days:
- The ETMY and ITMY pitch and yaw dofs are modulated at 40, 44, 42, 46 Hz respectively (oscillator A=30). The c1ass lockin numbers are 12, 14, 27, 29.
- The NAS55I signal is demodulated at the above frequencies. The demodulated I/Q signal phase is set to shift all signal into I-phase. The lockin inputs are bandpassed around respective frequency f with butter("Bandpass",2,f-0.5,f+0.5). The demod signals are then additionally low passed with butter ("Lowpass",4,0.5) so the servo ugf has to be below 0.5 Hz. The servo filter is p:z 0.0001:0.1.
- The ETMY demodulated signal is fed back to ITMY and visa versa.
- With the above 2x2 servo running we moved the input beam PZTs by hand to follow the cavity.
- At the end we offloaded the servo control signals to the SUS biases again by hand.
- The beam spot centering was estimated by unbalancing the ETMY/ITMY pitch/yaw coil combinations intentionally by 5%, which produces 1.3 mm shift of the node, and comparing the response to the residual signals.
- The dof set up currently is: ETMY pitch lockin 12 -> dof2, ITMY pitch lockin 14 -> dof4, ETMY yaw lockin 27 -> dof7, ITMY yaw lockin 29 -> dof9
- The next step is to demodulate the TRY(X) and servo the input beam PZTs
I put the ETMY trans QPD in.
The ETMY trans beam was already going toward the TRY DC PD, and a CCD camera. I put a beam splitter in that beam (reducing the power to TRY and the CCD by 50%), and sent my picked-off beam to the ETMY QPD. Since there is a lens in this path to focus the beam onto TRY and the camera, I put the QPD ~the same distance from the lens as the camera. Due to space requirements (because of all the green stuff on the table now), I had to put a Y1 turning mirror between the beam splitter and the QPD. The beam is aligned onto the PD, although the signal isn't super strong. When the PD is blocked, the sum is ~(-92 counts). When the beam is on the PD, the sum is ~(-78 counts).
Rana forced me to write this entry for summary because he didn't come to the 40m meeting today.
Status update :
Interferometer Input Beam alignment with the PZTs.
60 % done. The rest of the 40 % is to make the procedures automated.
The beam spots on ITMY and ETMY are centered within ~1 mm accuracy.
PRM, BS, ITMX, & ITMY actuator calibrations and PRC/MICH error calibrations
Ongoing: First we will do it by hand, then some scripts will be made for the calibraion and resultant noise budget.
F2A Suspension filter calculations.
ETMY and ITMY are done. Need volunteer for ETMX, ITMX, BS, PRM, & SRM !!
Bounce-Roll notch filters
Leo is working on it. 25% complete...
DC signal from RFPDs
The RFPDs have a local SMA DC output as well as a DC output from their PD Interface cards in the LSC rack. We have hooked up some of the PD Interface DC outputs to the LSC ADCs but not tested??
Installation of a temporary (Thorlabs) DCPD on either POY to see the intra-cavity power in the PRC. It would be ridiculous to put detectors on POX or POP since they're still clipped.
D-phase and amplitude imbalance adjustment of the demod baords.
make a script which uses pynds.
Alignment of the full interferomter, starting from the X arm
Loss measurements for the arms
Schnupp asymmetry measurement
I found that a He-Ne laser which has been used for ETMX_OPLEV was NOT giving the light.
Since I didn't find the switch key for it I have no idea if the laser is simply off or dead.
The dead laser was replaced by new JDSU 1104P of 2.6mW. The return beam is big ~5 mm diameter of 0.3 mW, 1400 counts
Whenever replacing any Oplev laser, please also put into the ELOG when it was installed so that we have an electronic record of the laser lifetime.
[Valera / Kiwamu]
The pointing of the incident beam to the interferometer has been jumping frequently.
Due to this jump the lock of the Y arm didn't stay for more than 2 min.
We turned off the strain gauge loop of PZT2-YAW and PZT2-PITCH, then the spot motion became solid and the Y arm locking became much more robust.
It was because of a loose connection. Pushing the connector solved the issue.
We really have to think about making reliable connections and strain reliefs.
Yesterday we found that MC3 OSEM LL PD did not have a sensible signal - the readback was close to zero and it was making MC move around. I disabled the PD LL so that the damping is done with just three face plus side PDs. There still no signal from MC3 LL PD today. It needs debugging.
The schematic for the seismometer box from this last time has been updated...
Koji was helpful for coming up with a general diagram for the voltage buffer amplifier, which has now been added to the configuration pictured below.
The only thing that remains now before I try to plot it with Eagle/LISO is to pick an op amp to use for the voltage buffer itself. Someone suggested THS4131 for that (upon Googling, it hit as a "high speed, low noise, fully-differential I/O amplifier"). It looks good, but is it the best option?
Aux- rack _1Y2 is just behind 1Y3 It contains Kepco Analoge DC power supplies for +- 5, 15 and 24V
Placing these power supplies away from the LSC rack was an effort to minimize pick up from them.
New f2p filters were installed on ITMY this morning. The statistical error of the coil gain setttigs are about 0.6 % at high frequency.
NEXT : PRM, SRM, BS, ITMX and ETMX.
Pendulum mode = 0.988 Hz, Q = 1
C1:SUS-ITMY_ULCOIL_GAIN = 1.02482
C1:SUS-ITMY_URCOIL_GAIN = -1.06831
C1:SUS-ITMY_LLCOIL_GAIN = -0.996671
C1:SUS-ITMY_LRCOIL_GAIN = 0.91079
UL: fz = 1.014824 Hz, Q = 1.027150
UR: fz = 0.975038 Hz, Q = 0.98688
LL: fz = 1.000229 Hz, Q = 1.012378
LR: fz = 0.972688 Hz, Q = 0.946116
New f2p filters were installed on ETMY.
The coil imbalances are now about 0.8% at high frequency (i.e. above the resonant freq of the pendulum mode)
The SRM qpd cable was removed from the BS-table. It's path was changed from 1x4 to ITMY-table following the inner cable tray.
Laser diode oplev SRM is working. Qpd matrix values were reset like others.
In 0.44mW, returning 0.1mW, -500 counts.
The dead laser was replaced by new JDSU 1103P of 2.6mW. The return beam is big ~5 mm diameter of 0.3 mW, 1400 counts
[Jamie / Valera / Kiwamu]
The incident beam pointing was improved a lot by using C1ASS realtime code.
Some more details will be posted later. The below is the list of the highlights today.
- The Y arm cavity was aligned to have good beam centering on the mirrors.
- The input PZTs were also aligned to the aligned Y arm by hand.
- Automation of the Y arm alignment using C1ASS_LOCKIN got partially functional with two loops closed. C1ASS correctly servos the centering on ETMY
- The amount of the off-centering on ITMY and ETMY look roughly within 1 mm.
- As a result the intracavity power got bigger by a factor of about 3.5
The statistical error of the coil gain settings are now about 0.8% at high frequency (i.e. above the resonant freq of the pendulum mode)
What I did :
- measured and corrected the coil imbalances on ETMY using a script called F2P_LOCKIN.py
- made the new f2p filters based on the measurements and installed them.
Next step :
- do the same adjustment for all the suspensions including PRM, SRM, BS, ITMs and ETMs
(Notes on F2P_LOCKIN.py)
F2P_LOCKIN.py is a script that I've made in python. This is basically the same as the old script, f2pratio, but uses the realtime LOCKINs instead of ezcademods.
The script automatically measures the coil imbalances on an optics of interest by driving the local LOCKIN oscillators.
In the first step the script automatically balances the coil gains at high frequency (8.5Hz).
In the next step it gives some coefficients, which basically represent the coil imbalances at low frequency (0.01Hz)
Then with those coefficients one will be able to design the f2p filters.
It is not well polished yet, so I will spend some more times to make it user-friendly and readable.
Example usage : F2P_LCKIN.py -o ETMX
It currently resides in /cvs/cds/rtcds/caltech/c1/scriptss/SUS/
(new f2p filters)
The plot below shows the new f2p filters. Note that they are already installed.
I re-installed the box (@ ~8:15) after reflowing some of the solder joints. I will observe it over night and then remove the 1K resistors. Attached is a 8 hour minute-trend.
I compared this 24 hour trend with the one from this day exactly one year ago. Seems the same, so now I can make the resistor change.
We added a cdsPCIx_SHMEM connection between the c1lsc and c1ass models. This connection is called C1:LSC-ASS_AS55I, and sends the normalized AS55I data to Lockin 11 of the c1ass model.
In addition, in order to get the c1ass model to compile, we had to place all the non-IO parts inside a subsystem block, which we called ASS, and gave the top_names tag to.
The c1lsc and c1ass models were rebuilt, the frame builder restarted, and the models restarted.
Ryan added the c1uct (upconversion tester) model to the c1ioo machine. It is DCU_ID 32, CPU 6. The relevant wiki page has been updated. It has been added to /diskless/root/etc/rtsystab on the fb machine and automatically comes up when the c1ioo computer is turned on.
Joe still needs to add its status to the status screen.
It is soft linked from:
Ryan will expand upon its uses later.
Valera and I placed F 572.7 mm lens ~15 cm away from the ang qpd (in the same mount with ND filter) so that two qpds see different combination of ang and pos motion - there was no lenses prior to this change. The beam diameter is reduced to ~half .
I added calculation entries to the /cvs/cds/caltech/target/c1iool0/c1ioo.db file which are named C1:IOO-QPD_*_*, as the channels were originally named. These calculation channels have the identical data to the C1:PSL-QPD_*_* channels. I then added the channels to the C0EDCU.ini file, so as to once again have continuity for the channels, in addition to having the newer, better named channels.
The c1iool0 machine ("telnet c1iool0", "reboot") and the framebuilder process ("telnet fb 8087", "shutdown") were both restarted after these changes.
These channels were brought up in dataviewer and compared. The approriate channels were identical.
First we changed all the C1:IOO-QPD_*_* channels to C1:PSL-QPD_*_* channels in the /cvs/cds/caltech/target/c1iool0/c1ioo.db file, as well as the /opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini file. We then rebooted the frame builder via "telnet fb 8087" and then "shutdown".
This change breaks continuity for these channels prior to today.
- The BBPD circuit has been constructed on the aLIGO BBPD board
- It still keeps 200MHz BW with FDD-100 Si PD for the 100V bias.
- The noise spectrum has been cleaned up a lot more. It shows the noise level of the 0.4mA shotnoise between 9-85MHz.
The noise at 160MHz is the noise level of the 1mA shotnoise.
Some of the noise peaks at around 97MHz came from the bias voltage.
What to do next
- Confirmation of the performance with the original aLIGO BB PD configuration.
- Notch filter for 9MHz (for aLIGO).
- Implementation of a power amplifier. (issues: power supply and heat removal)
Today after Steve and I finished the RF cabling work for the day, Kiwamu noticed that there were no RF signals to be seen. The problem was traced to disconnected 11 and 55 MHz Demod lines from the RF source. But reconnecting them did not restore the signals. It turned out that one of the Heliax cables had a loose N-type connector at its end and it finally came off while we were tightening it into place.
We replaced the damaged heliax with another (we have two spare running from 1X2 IOO rack to the 1Y2 LSC rack. The new cable is used to be the LO 33. It seems to have a 1.5dB loss. Have to check this again tomorrow.
In the mean time I noticed that the power output of the 55MHz Demod port of the source was less than about -12dB. So I opened the source to take a look and found that all the voltage stabilisers were supplying 15V. Even those which were supposed to be supplying 24V. This was traced to a mistake in wiring the external power supply. The wires had been labeled wrongly and as a result the 18V input line was connected to 28V source and vice versa.
After fixing this problem I reassembled the source checked the power output on all the ports and found everything was functioning as expected. However after installation once again the unit failed. The blue light on the power supply was not lighting up when switched on. Suspecting a power supply problem I opened the unit again and found that a weak solder joint on one of the RF amplifiers had come loose and had overloaded one of the 24V stabilisers. We, found a spare and replaced it. The unit has been reassembled and is functioning fine. The output power levels are
11MHz Demod -- 6dBm
55MHz Demod -- 5.5 dBm
11MHz EOM -- 24dBm
55MHz EOM -- 28dBm
The Marconi is serving as the 11MHz source. The Wenzel 11MHz source is giving 13.3 dBm and is okay. But it needs to be checked for its performance as it may have been exposed to higher than rated power supply levels.
The 29.5MHz source is giving 7dBm. It is supposed to be giving 13dBm.
The Laboratory DC power supplies currently used for both the RF source and Distribution boxes need to be replaced with rack mounted Sorensen power supplies available in the lab.
Kiwamu, Koji, Valera
We centered the beam on MC2 in pitch by moving the MC1,2,3 in the following combination [-9,+3,-7]. This actuation vector mostly moves the spot on MC2 vertically. The attached plot shows the dither before and after the centering. We monitored the demodulated signals and saw the reduction of the MC2 pit response from -1.0 to -0.22 which corresponds to the beam spot position change from 9 to 2 mm. Thus all the spots on MC mirrors are within 2 mm of the center. We estimate based on the distance between the MC1-MC3 of 20 cm, the distance from the center between MC1 and MC3 to the end of the Faraday isolator of 80 cm, and the aperture of the FI of 12 mm, the maximum angle out of MC of 3/200 rad. Which implies the maximum differential spot motion of 3 mm not to be limited by the FI aperture.
The alignment of the interferometer goes basically step by step.
Tuesday will be an alignment day.
0. MC beam centering (it's done)
1. F2P to balance the coils on every optics including BS, PRM, SRM, ITMs and ETMs (Kiwamu).
2. A2L and then change the DC bias of ITMY and ETMY to get a perfect eigen axis (VF/Jamie).
3. align input PZT mirrors (PZT1 and 2) to maximize the Y arm transmission (VF/Jamie).
4. do the same things for X arm but using BS instead of the PZTs.
5. Alignment of the central part.
6. Make a script to automatically get those things done.
I made a new medm screen for the triggering logics. Have fun.
I put a button on C1LSC.adl to invoke this screen.
It will be quite useful for the Y arm locking, for instance we can do a triggered locking and the maximization of the intracavity power.
We started to clean up the RF cables (heliax and PD interface cables) at the LSC rack.
We have pulled out all the RF cables from the small hole on the side-board close to floor. Passing the cables through this hole makes some of the cables much too short for good strain relief. So we removed the side panel on the vacuum tube side and are going to pass the cables into the rack from there at about waist height. We now have plenty of cable lengths to tie them off to the rack at several points.
We have traced all the available Heliax cables and have attached blank tags to them. We have allocated some cables to REFL11, REFL55 and AS55. These are therefore back in working order. We have also taken stock of the available PD interface cables. They do not have consistent names on both ends of the cable and we will identify and label the ends tomorrow.
MC is locked. The auto-locker works fine.
Handing over the system for night time interferometer work now. Will continue with the cabling tomorrow.
Also, remember to update the svn working copy before you begin doing any work, to make sure you're working on the most recent version:
Whenever you modify any of the front-end code in /opt/rtds/caltech/c1/userapps, such as models, scripts or medm screens, REMEMBER TO COMMIT YOUR CHANGES, WITH A LOG MESSAGE!!!
svn commit path/to/modified/file
If you forget to commit things, they may be purged and you will loose your work. ALWAYS COMMIT!
Things to note and watch out for:
If you have questions ASK. Don't force things.
I have updated the /opt/rtcds paths to reflect the new specification of the CDS aLIGO code release procedures document.
Path to RTS: /opt/rtcds/caltech/c1/core/release
This is where the advLigoRTS front-end code generator is checked out. The "release" directory is a link to the svn branch from which we are currently running ("trunk" by default).
This used to be at /opt/rtcds/caltech/c1/core/advLigoRTS
Path to userapps: /opt/rtcds/caltech/c1/userapps/release
This is where the cds_user_apps code is checked out. cds_user_apps is where all of the front-end models, medm screen, scripts, etc. will live. The "release" directory is a link to the svn branch from which we are currently running ("trunk" by default).
This was most recently at /opt/rtcds/cds_user_apps
The C1:PSL-QPD_POS_HOR and C1:PSL-QPD_POS_VERT channels were found to be backwards as well. So we modified the /cvs/cds/caltech/target/c1iool0/c1ioo.db file to switch them.
Lastly, we changed the ASLO and AOFF values for the C1:PSL-QPD_POS_SUM and the C1:PSL-QPD_ANG_SUM so as to provide positive numbers. This was done by flipping the sign for each entry.
ASLO went from 0.004883 to -0.004883, and AOFF when from -10 to 10 for both channels.
The C1:PSL-QPD_ANG_SUM channel had been saturated at -10V. Valera reduced the power on the QPD to drop it to about 4V by placing an ND attenuator in the ANG QPD path.
There is a feature/bug of the RCG code that you can only have 1 receiving tag for every sending tag. There were 5 tags which were being received by two tags each, for two different matrices. Only the first tag was receiving, the second was apparently ignored.
This has been fixed temporarily by putting in direct lines in place of these 5 tags.
I found that C1:LSC-TRIG_MTRX has a wrong matrix size. It needs to be fixed.
It is designed to have a 11x8 matrix in the simlink model file, but it's been compiled as a 6x8 matrix.
I recompiled c1lsc but it didn't fix the issue. Here below is the matrix statement in the C file :c1lsc.c. Indeed it is 6x8 matrix ....
// MuxMatrix: LSC_TRIG_MTRX
pLocalEpics->lsc.LSC_TRIG_MTRX[ii] * lsc_imux_trigger +
pLocalEpics->lsc.LSC_TRIG_MTRX[ii] * lsc_imux_trigger +
pLocalEpics->lsc.LSC_TRIG_MTRX[ii] * lsc_imux_trigger +
pLocalEpics->lsc.LSC_TRIG_MTRX[ii] * lsc_imux_trigger +
pLocalEpics->lsc.LSC_TRIG_MTRX[ii] * lsc_imux_trigger +
pLocalEpics->lsc.LSC_TRIG_MTRX[ii] * lsc_imux_trigger;
The attached plot shows 7 day trends of the MC and PMC power levels, PSL QPDs, and temperature. The MC stayed locked for ~40 hours over the weekend. The temperature swings were somewhat smaller over the past couple of days but one should remember to turn the PSL HEPA down after working on the table. Steve turned the HEPA flow from 100% down to 20% on Thursday and posted the reminder signs on the PSL enclosure.
I looked at the PSL temperature box. It started out as D980400-B-C. Then it was revised by Peter King as per the LHO mods E020247.
There are some more things to do to it to make it useful for us:
** Frank reminds me that we don't use the TIdal or VME external inputs anymore since we moved to the EPICS/Perl PID control. So all we have to do is make sure these inputs are hardware disabled/disconnected.
It seems like the best option would be to make the MCASS just adjust the SUS biases and center the beams on the suspended optics. Is this not possible somehow?
Kiwamu told me that the CDS matrix notation has changed and the 40m front end code has changed since February. I changed the senseMCdecentering script to reflect that. The other problems were: the "-" sign in ezcastep on ubuntu is not recognized - I used the known workaround of using "+-" instead; the echo command in csh script on ubuntu does not make a new line - but the echo " " does. The script ran on ubuntu with one error message "FATAL: exception not rethrown" but it finished nevertheless. The data appeared ok. On centos machine the script produced "Segmentation fault'. The matlab script sensemcass.m now calculates the position on the MC mirrors in mm. The attached table shows the MC spot positions in mm:
I had to rephase the lockin digital phases by tens of degrees. I don't know why this should happen at ~10 Hz.
I tried to run the scripts/senseMCdecentering to check the centering of the MC beam spots on the mirrors. The script (csh) produces a lot of error messages on the control room machines. They are machine dependent combination of "epicsThreadOnce0sd epicsMutexLock failed", "Segmentation fault", "FATAL: exception not rethrown". Most of ezcawrite commands fail but not all(?). After running the mcassUp script couple of times all the dither lines came on. The MCL responses to dither lines look qualitatively similar to what it was in February (plot attached). The overall MCL spectrum looks ~100 times lower, presumably due to the analog gain reallocation.
Before that I realigned the beam into the PMC, recentered the PSL QPDs, and the beam into the MC to bring the MC RFPD_DC from ~3 to ~1.5 VDC then tweaked MC2 to bring the MC RFPD_DC from ~1.5 to ~1 VDC.
The mcass dither lines are off now and the loops are disabled.
The RF amplifier of the prototype BBPD has been replaced from ERA-5SM to MAR-6SM.
The bandwidth is kept (~200MHz for S3399 with 30V_bias), and the noise level got better while the maximum handling power was reduced.
MAR-6SM is a monolithic amplifier from Minicircuits. It is similar to ERA-5SM but has lower noise
and the lower output power.
The noise floor corresponds to the shotnoise of the 0.4mA DC current.
Now the mess below 50MHz and between 90-110MHz should be cleaned up.
They are consistently present no matter how I change the PD/RF amp (ERA<->MAR)/bias voltage.
I should test the circuit with a different board and enhanced power/bias supply bypassing.
- Assume 5mA is the maximum RF (~50mW for 1064nm, ~15mW for 532nm). This is already plenty in terms of the amount of the light.
- 100% intenisty modulation for 5mA across 50Ohm induces -2dBm RF power input for the amplifier.
- Assume if we use MAR-6 for the preamplifier. The max input power is about -18dBm.
This corresponds to 16% intensity modulation. It may be OK, if we have too strong intensity modulation, we can limit the power
down to 0.8mA in the worst case. The shot noise will still be above the noise level.
- In the most of the applications, the RF power is rather small. (i.e. 40m green beat note would expected to be -31dBm on the RF amp input at the higherst, -50dBm in practice)
So probably we need more gain. If we can add 10-12dB more gain, that would be useful.
- What is the requirement for the power amplifier?
Search result for Freq Range 10-200MHz / Max Gain 14dB / Max NF 15dB / Min Power Out 13dBm
GVA-81 is available at the 40m. ERA-4SM, ERA-6SM, HELA-10D are available at Downs.
Conversion between nV/rtHz and NF (in the 50Ohm system)
SN1: Connect signal source (50Ohm output) to a 50Ohm load.
Power ratio between the noise and the signal
SN1 = (4 k T (R/2)) / (S/2)^2
SN2: Connect signal source (50Ohm output) to an RF amp.
Only the voltage noise was considered.
SN2 = (4 k T (R/2) + Vn^2) / (S/2)^2
10 Log10(SN2/SN1) = 10. Log10(1 + 2.42 (Vn / 1nVrtHz)^2)
Vn: 0 nVrtHz ==> 0dB
Vn: 0.5 nVrtHz ==> 2dB
Vn: 1 nVrtHz ==> 5dB
Vn: 2 nVrtHz ==> 10dB
Vn: 3 nVrtHz ==> 13.5dB
The SLP-50 filters which were on the 55 MHz lines have been replaced with the SBP-60. Their respective characteristics are given below:
SBP-60 has lower insertion loss and higher return loss.
This may however change the phase of I and Q in the demod boards and they will therefore need to be readjusted. Currently the output power level of 55 MHz demod is at 2dBm, whereas it ought to be at 6dBm. I have not yet corrected that. Once that is completed Kiwamu will adjust the phases.
I shifted the temperature sensor to a new location. See the photograph below. I noticed that the higher temperature is reached on the side where there are two RF Amps. So it would be better to check the temperature of that area and make sure that it remains well below 65 deg. The operating maxium is 65deg C
Here is a picture of the new RF source layout.
And here is a photograph of it
[Suresh / Kiwamu]
We installed the TRY photo diode (Thorlabs one) and the ETMYT CCD camera in place on the ETMY table.
Now we can see a signal on the TRY digital channel.
Someone has to install the EMTY trans QPD at some point.
Dry fore pump of TP3 replaced by brand new Varian SH-110 at 1.1 Torr_75,208 hrs
The annuloses were closed off for 25 minutes. We are back to VACUUM NORMAL mode.
The TP3 fore line pressure dropped to 44 mTorr at 25 minutes in operation.......9.4mTorr at day 2 with full annulos load
After Kiwamu set the REFL11 phases in the PRMI configuration (maximized PRM->REFL11I reesponse) I tried to measure the MC length and the 11 MHz frequency missmatch by modulating the 11 MHz frequency and measuring the PM to AM conversion after the MC using the REFL11Q signal. The modulation appears in the REFL11Q with a good snr but the amplitude does not seem to go through a clear minimum as the 11 MHz goes through the MC resonance.
We could not relock the PRMI during the day so I resorted to a weaker method - measuring the amplitude of the 11 MHz sideband in the MC reflection (RF PD mon output on the demod board) with a RF spectrum analyzer. The minimum frequency on the IFR is 11.065650 MHz while the nominal setting was 11.065000 MHz. The sensitivity of this method is about 50 Hz.
I tuned the ITMY bandstops -- 'before' and 'after' spectra attached. Note that the after the tuning, the bounce mode at ~16 Hz is about twice as quiet!
However, notice that in the 'before' plot the roll mode at about 23.5 Hz did not show up at all, whereas it is quite prominent in the 'after' plot. I was concerned that this line could have been a result of placing the bandstop there, so I made another plot with the BounceRoll filter turned off. Sure enough, the 23.5 Hz line is still there. So I'm not crazy: the roll mode did start acting up at some time between my 'before' and 'after' plot, but not as a result of the tuning.
The emphasis of this annual safety audit was on safe electrical housekeeping on March 3, 2011
We now have the DC signal from three PDs available in the ADC channels 14,15 and 16. The signals are from REFL55, AS55 and POY photodiodes respectively. As the DC signals on all the other PDs of the same port (REFL, AS and PO) have the same information we do not need to monitor more than one DC PD at each port.
The LSC PD Interface Card, D990543 - Rev B, can take 4 PDs and provides the DC signals of the PDs on the connector P2 (the lower of the two) on the back plane of the chassis. An adaptor card, D010005-00, plugs into the back plane from the rear of the Eurorack and provides the four DC signals on two-pin lemo sockets.
I have connected the three DC signals from the relevant RF PDs (above) to a DC whitening filter, D990694-B-1 which is associated with the channels 9 to 16 of the ADC card.
The cables are in a bit of a mess right now as some of the PD power supply lines are too short to reach up the the Interface card in the top Eurocart. Steve and I plan to redo some of the cabling later today
Minicircuits ERA-5SM was used for the RF amp of the BBPD. This amp is promising as a replacement of Teledyne Cougar AP389
as ERA-5SM gave us the best performance so far among the BBPDs I have ever tested for the aLIGO BBPD/Green.
The -3dB bandwidth of ~200MHz and the noise floor at the shotnoise level of 0.7mA DC current were obtained.
The aLIGO BBPD candidate (LIGO Document D1002969-v7) employs Teledyne Cougar AP389 as an RF amplifier.
This PD design utilizes the 50Ohm termination of the RF amp as a transimpedance resistance at RF freq.
However, it turned out that the bandwidth of the transimpedance gets rather low when we use AP389, as seen in the attachment2.
The amplifier itself is broadband upto 250MHz (the transfer function was confirmed with 50Ohm source).
The reason is not understood but AP389 seems dislike current source. Rich suggested use of S-parameter measurement
to construct better model of the curcuit.
On the other hand, the RF amplifiers from Minicircuits (coaxial type like ZFL-1000LN+), in general, exhibit better compatibility with PDs.
If you open the amplifier case, you find ERA or MAR type monolithic amplifiers are used.
So the question is if we can replace AP389 by any of ERA or MAR.
- The large gain of the RF amp is preffered as far as the output does not get saturated.
- The amplifier should be low noise so that we can detect shot noise (~1mA).
- The freq range of the useful signal is from 9MHz to 160MHz.
The advanced LIGO BBPD is supposed to be able to receive 50mW of IR or 15mW of 532nm. This approximately corresponds to
5mA of DC photocurrent if we assume FFD-100 for the photodiode. At the best (or worst) case, this 5mA has 100% intensity modulation.
If this current is converted to the votage through the 50Ohm input termination of the RF amp, we receive -2dBm of RF signal at maximum.
This gives us a dilemma. if the amp is low noise but the maximum output power is small, we can not put large amount of light
on the PD. If the amp has a high max output power (and a high gain), but the amp is not low noise, the PD has narrow power range
where we can observe the shotnoise above the electronics noise.
What we need is powerful, high gain, and low noise RF amplifier!
Teledyne Cougar AP389 was almost an ideal candidate before it shows unideal behavior with the PD.
Among Minicircuits ERA and MAR series, ERA-5 (or ERA-5SM) is the most compatible amplifier.
Considering the difference of the gain, they are quite similar for our purpose. Both can handle upto -2dBm,
which is just the right amount for the possible maximum power we get from the 5mA of photocurrent.
A test circuit has been built (p.1 attachment #1) on a single sided prototype board.
First, the transfer function was measured with FFD-100. With the bias 100V (max) the -3dB bandwidth of ~200MHz was observed.
This decreases down to 75MHz if the bias is 25V, which is the voltage supplied by the aLIGO BBPD circuit. The transimpedance
at the plateau was ~400Ohm.
Next, S3399 was tested with the circuit. With the bias 25V and 30V (max) the -3dB bandwidth of ~200MHz was obtained although
the responsivity of S3399 (i.e. A/W) at 1064nm is about factor of 2 smaller than that of FFD-100.
The noise levels were measured. There are many sprious peaks (possibly by unideal hand made board and insufficient power supply bypassing?).
Othewise, the floor level shows 0.7mA shotnoise level.